A certain rifle bullet has a mass of 8.85g.  Calculate the de Broglie wavelength of the bullet traveling at 1649 miles per hour. Consult the table of physical constants as needed.
Answer #1

1 miles = 1.609 km = 1609 m

Given:

v = 1649 miles / hour

= 1649 * 1609 m / 3600 s

= 737.0 m/s

m = 8.85 g

= 8.85*10^-3 Kg

use:

lambda = h / (m*v)

lambda = (6.626*10^-34) / (8.85*10^-3 * 7.37*10^2)

lambda = 1.016*10^-34 m

Answer: 1.016*10^-34 m